Alpha Company, on March 1, 2017 has a beginning Work in Process inventory of zero. All materials are added into production at the beginning of its production. There is only one production WIP inventory. On March 1, Alpha started into production 10,000 units. At the end of the month there were 5,000 units completed and transferred into the Finished Goods Inventory. The ending WIP was 60% complete with respect to conversion. For the month of March the following costs were incurred and recorded in the WIP: Direct Material $20,000 Direct Labor 8,000 Factory Overhead 32,000 Alpha uses the weighted-average process costing method. Use this information to determine for the month of March: (Round & enter final answers to the nearest penny.)

1. the cost per equivalent unit of direct material

2. the cost per equivalent unit of conversion

3. the cost per unit transferred to finished goods
Answer #1

1) Equivalent unit of material = 10000

Cost per equivalent unit = 20000/10000 = 2

2) equivalent unit of conversion = (5000+5000*60%) = 8000

Cost per equivalent unit of conversion = (8000+32000)/8000 = 5

3) Total cost per unit transferred to finished goods = 5+2 = 7 per unit